Circular No. 5925

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 S. N. Shore, Indiana University, South Bend; S. Starrfield and P. H.

Hauschildt, Arizona State University; G. Sonneborn, Goddard Space Flight Center; and R. González-Riestra, International Ultraviolet Explorer (IUE) Observatory, Vilspa, report: “Observations of N Cas 1993 have been obtained with the IUE satellite since shortly after the report of its discovery (IAUC 5902). The first spectra (1993 Dec. 12.1 UT) showed the nova at ultraviolet (UV) minimum with most of the opacity coming from overlapping absorption lines from the iron group elements: the ‘iron curtain’. We have been obtaining IUE spectra at ∼ 4-day intervals since that time, which reveal that the ejected material has remained in an extremely optically thick shell stage. The wavelength region around 280 nm is a blend of both Fe II and Mg II. This blend displays a pseudo-P-Cyg profile with a broad absorption trough that extends to a velocity of ∼ −1600 km/s. Interstellar Mg II h and k absorption lines are superposed on emission from the nova; they have mean velocities of −35 km/s but extend to nearly −60 km/s. Windows in the Fe II absorption spectrum have recently begun to appear at 158 nm and longer wavelengths, and changes in the spectral energy distribution imply a slowly growing temperature for the ejecta. An ultraviolet rise of a factor of three occurred around Dec. 22, but after that time the nova has remained at nearly constant integrated UV flux (through 1994 Jan. 12). A preliminary upper limit of E(B–V) = 0.6 is derived from the UV-to-optical flux ratio. The interstellar Mg II and Na I lines suggest a distance of 4-6 kpc. We note that the optical light curve is similar to that of a DQ Her-type of nova (Payne-Gaposchkin 1957, The Galactic Novae, p. 13). Such a nova will form an optically-thick dust shell ∼ 60-90 days after maximum light in the optical. Our stellar-atmosphere modeling indicates that the slow spectral evolution is consistent with steady-state ejection over the past month. The optical light curve, the relatively slow UV spectral evolution, the steady state mass ejection, and the detection of CO molecular lines in the infrared (IAUC 5916, 5922) lead us to expect that this nova will also form optically-thick dust (Shore and Starrfield 1994, Sky Telesc., in press); infrared observations will be necessary to confirm this expectation.”

H. Mikuž, Ljubljana, Slovenia, reports that a prediscovery photograph, taken by J. Prudic (Ljubljana) on 1993 Dec. 4.95 UT with a 58-mm f/2 lens and T-Max 400 film, shows this object at m_{pv} = 8.0 (comparison stars from AAVSO Atlas).

1994 January 20 Daniel W. E. Green